# Multimodal 3D Reasoning Segmentation with Complex Scenes

### v1

@@ Abstract
The recent development in multimodal learning has greatly advanced the research in 3D scene understanding in various real-world tasks such as embodied AI. However, most existing work shares two typical constraints: 1) they are short of reasoning ability for interaction and interpretation of human intension and 2) they focus on scenarios with single-category objects only which leads to over-simplified textual descriptions due to the negligence of multi-object scenarios and spatial relations among objects. We bridge the research gaps by proposing a 3D reasoning segmentation task for multiple objects in scenes. The task allows producing 3D segmentation masks and detailed textual explanations as enriched by 3D spatial relations among objects. To this end, we create ReasonSeg3D, a large-scale and high-quality benchmark that integrates 3D spatial relations with generated question-answer pairs and 3D segmentation masks. In addition, we design MORE3D, a simple yet effective method that enables multi-object 3D reasoning segmentation with user questions and textual outputs. Extensive experiments show that MORE3D excels in reasoning and segmenting complex multi-object 3D scenes, and the created ReasonSeg3D offers a valuable platform for future exploration of 3D reasoning segmentation. The dataset and code will be released.
@@ 1 Introduction
On top of ReasonSeg3D, we design MORE3D, a simple yet effective technique that enables multi-object reasoning segmentation with textual explanations to users’ questions. MORE3D learns object-specific point cloud embeddings from LLM, enabling predicting precise 3D segmentation masks of multiple objects. In addition, it produces detailed textual explanations that capture 3D spatial relations among multiple objects, supporting accurate segmentation and comprehensive reasoning for complex 3D scenes. As illustrated in Figure 1, MORE3D demonstrates strong reasoning capability for comprehending the intention behind user’s input questions and complex 3D scenes, producing accurate 3D segmentation and explanatory answers with respect to multiple 3D objects in scenes.
@@ 3.2 Dataset Generation Pipeline
We design a novel data generation pipeline that introduces GPT-4o that possesses superior visual content understanding capabilities. The pipeline allows generating more practical questions and answers by incorporating spatial relations among objects in scenes. Specifically, we employ both scene image and its ground-truth segmentation to enhance GPT-4o’s 3D spatial understanding capabilities. The prompt template is structured in two parts: the first part contains basic requirements for the GPT-4o and the second specifies detailed requirements for generating questions and answers with a focus on describing 3D spatial relations among objects. With this template, GPT-4o autonomously selects objects to form question-answer pairs that reflect the scene’s content and spatial layout. Figure 2 shows one example of the designed prompt template and one question-answer pair generated by using the template.
@@ 4.3 Multi-Object 3D Reasoning Segmentation
We design an object-specific point embedding extraction approach to achieve the multi-object 3D reasoning segmentation. The embedding extraction approach enables precise segmentation and reasoning for multiple objects of different categories in complex point clouds, integrating the LLM with a 3D point cloud decoder to generate segmentation masks and textual explanations.
@@
Figure 5: Segmentation visualization over the ReasonSeg3D validation set. Each case presents a user input question, the corresponding input point cloud, the ground-truth segmentation, and the prediction by the proposed MORE3D. Best viewed in color and zoom-in.
@@ 5.2 Benchmarking with Existing Methods
We benchmark MORE3D with state-of-the-art 3D segmentation methods on the ReasonSeg3D validation set. As Table 2 shows, our method achieves superior segmentation performance across all evaluation metrics. Specifically, OpenScene [30], PLA [9], and RegionPLC [45] are designed to comprehend input words or phrases based on CLIP-based language models. They lack the ability to reason user intention from input questions and to correctly generate answers targeting at corresponding multiple objects. Consequently, the predicted 3D segmentation masks by these methods cannot be aligned correctly with ground-truth 3D segmentation masks, resulting in low cIoU and gIoU. In contrast, MORE3D can accurately reason the implicit intention of user questions, predicting 3D segmentation masks that are more consistent with the ground truth.
@@ 5.3 Ablation Study
We examine how different point cloud decoding operations affect 3D reasoning segmentation. Specifically, we evaluate three approaches of point cloud decoding operations that integrate per-point features and object embeddings to predict the 3D segmentation masks. As Table 5 shows, the Addition and the Concatenation operations achieve lower performance compared with the Dot Product operation since they either add or concatenate the embeddings of each object with the per-point features of the global scene, introducing abundant features of irrelevant objects and leading to inferior performance. Differently, the Dot Product operation selectively extracts only the relevant features from the per-point features of the global scene by identifying those highly correlated with the object embeddings. This leads to better performance thanks to its extracted object-specific features.

### v2

@@ Abstract
The recent development in multimodal learning has greatly advanced the research in 3D scene understanding in various real-world tasks such as embodied AI. However, most existing work shares two typical constraints: 1) they are short of reasoning ability for interaction and interpretation of human intension and 2) they focus on scenarios with single-category objects only which leads to over-simplified textual descriptions due to the negligence of multi-object scenarios and spatial relations among objects. We bridge the research gaps by proposing a 3D reasoning segmentation task for multiple objects in scenes. The task allows producing 3D segmentation masks and detailed textual explanations as enriched by 3D spatial relations among objects. To this end, we create ReasonSeg3D, a large-scale and high-quality benchmark that integrates 3D segmentation masks and 3D spatial relations with generated question-answer pairs. In addition, we design MORE3D, a novel 3D reasoning network that works with queries of multiple objects and tailored 3D scene understanding designs. MORE3D learns detailed explanations on 3D relations and employs them to capture spatial information of objects and reason textual outputs. Extensive experiments show that MORE3D excels in reasoning and segmenting complex multi-object 3D scenes, and the created ReasonSeg3D offers a valuable platform for future exploration of 3D reasoning segmentation. The dataset and code will be released.
@@ 1 Introduction
On top of ReasonSeg3D, we design MORE3D, a simple yet effective technique that enables multi-object reasoning segmentation with textual explanations to users’ questions. Specifically, MORE3D extracts object-specific point cloud embedding for precise prediction of multi-category objects in complex 3D scenes. It incorporates textual explanations with detailed 3D spatial relations of multiple objects in answers, guiding the model toward a comprehensive understanding of 3D scenes. The generated textual answers support accurate segmentation and comprehensive reasoning for complex 3D scenes. As illustrated in Figure 1, MORE3D demonstrates strong reasoning capability for comprehending the intention behind user’s input questions and complex 3D scenes, producing accurate 3D segmentation and explanatory answers with respect to multiple 3D objects in scenes.
@@ 3.2 Dataset Generation Pipeline
We design a novel data generation pipeline that introduces GPT-4o that possesses superior visual content understanding capabilities. The pipeline allows generating more practical questions and answers by incorporating spatial relations among objects in scenes. Specifically, we employ both scene image and its ground-truth segmentation to enhance GPT-4o’s 3D spatial understanding capabilities. The prompt template is structured in two parts: the first part contains basic requirements for the GPT-4o and the second specifies detailed requirements for generating questions and answers with a focus on describing 3D spatial relations among objects. With this template, GPT-4o autonomously selects objects to form question-answer pairs that reflect the scene’s content and spatial layout. On top of that, we conduct human verification to ensure the quality of the generated question-answer pairs. Figure 2 shows one example of the designed prompt template and one question-answer pair generated by using the template.
@@ 4.3 Multi-Object 3D Reasoning Segmentation
We design an object-specific point cloud embedding extraction approach to achieve the multi-object 3D reasoning segmentation. The embedding extraction approach enables precise segmentation and reasoning for multiple objects of different categories in complex point clouds, integrating the LLM with a 3D point cloud decoder to generate segmentation masks and textual explanations.
@@
Figure 5: Segmentation visualization over the ReasonSeg3D validation set. Each case presents a user input question, the corresponding input point cloud, the ground-truth segmentation, and the prediction by the proposed MORE3D. Best viewed in color and zoom-in. Green indicates chairs, and pink indicates tables. Best viewed in color.
@@ 5.1 Experimental Settings
3D-STMN [41] AAAI 24 19.37 20.70
RefMask3D [12] MM 24 20.53 22.81
MDIN [40] MM 24 23.04 23.49
3D Vision-Language Methods
3D-LLM [14] NeurIPS 23 21.36 22.49
3D-VisTA [53] ICCV 23 23.95 24.77
@@ 5.2 Benchmarking with Existing Methods
We benchmark MORE3D with state-of-the-art 3D segmentation and vision-language methods on the ReasonSeg3D validation set. As Table 2 shows, our method achieves superior reasoning segmentation performance across all evaluation metrics. Specifically, OpenScene [31], PLA [9], and RegionPLC [46] are designed to comprehend input words or phrases based on CLIP-based language models. They lack the ability to reason user intention from input questions and to correctly generate answers targeting at corresponding multiple objects. Consequently, the predicted 3D segmentation masks by these methods cannot be aligned correctly with ground-truth 3D segmentation masks, resulting in low cIoU and gIoU. Moreover, 3D-STMN [41] and RefMask3D [12] are tailored for single-object segmentation and therefore cannot handle multiple objects, while MDIN [40] exhibits weaker reasoning ability for interpreting implicit human intention. Additionally, 3D-LLM [14] and 3D-VisTA [53], designed for general 3D tasks, are re-implemented on our proposed 3D reasoning segmentation task for comparison. In contrast, MORE3D can accurately reason the implicit intention of user questions, predicting 3D segmentation masks that are more consistent with the ground truth.
@@ 5.3 Ablation Study
We examine how different point cloud decoding operations affect 3D reasoning segmentation. Besides the dot product for refined queries and features, we explore addition and concatenation to investigate whether richer point cloud information improves training. As Table 5 shows, the Addition and the Concatenation operations achieve lower performance compared with the Dot Product operation since they either add or concatenate the embeddings of each object with the per-point features of the global scene, introducing abundant features of irrelevant objects and leading to inferior performance. Differently, the Dot Product operation selectively extracts only the relevant features from the per-point features of the global scene by identifying those highly correlated with the object embeddings. This leads to better performance thanks to its extracted object-specific features.